OpenStudy (anonymous):

A baseball is hit with a speed of 30m/s at an angle of 42 degrees. It lands on the flat roof of a 14.m tall nearby buildining. If the ball was hit when it was 1.4 m above the ground, what horizontaldistance does it travel before it lands on the builiding?

OpenStudy (anonymous):

Have I set the problem up right so far? a/g = 9.8 m/s^2, Vy0=30m/s sin(42) theta = 42, y = -14,, Vx0= 20m/scos(42)

OpenStudy (anonymous):

oh and Vy0 and Vx0 are initial velocity components

OpenStudy (nottim):

Always, always drawa diagram first for these equations.

OpenStudy (nottim):